# Towards a complete scheme of cosmological neutrino self-interactions: Collision term for a wide range of mediator masses

### v1

@@ 1 Introduction
From the data analysis point of view, the main difference between these approximations relies on the cosmological scales being affected, with the mediator mass having an inverse correlation to the scale. The very light mediator and resonant regimes affect the largest scales, modifying the CMB. On the other hand, for the heavy mediator, small-scale data, especially the full shape of the matter power spectrum as seen in the galaxy and Lyman-$\alpha$ forest distributions [28, 76, 29, 102, 77] (in some scenarios) prefers a large non-null interaction coupling that should be observed in particle detectors unless one proposes or recalls a mechanism to avoid experimental bounds. Thus, there is an abnormal signal that future data may be able to explore [15, 74, 79, 80, 19, 22, 26, 14, 106, 57, 32, 27, 24, 47, 56, 109, 88, 38, 114, 110, 13, 41, 108, 60, 55, 31, 59, 68, 91, 1, 35, 85, 7, 33, 61, 53, 30, 51, 105, 63, 62, 23, 6, 92, 49, 46, 120, 43, 66, 115, 40, 67, 2].
@@ 2.1 Boltzmann hierarchy equations under cosmological perturbations
It is clear that $Q_{\alpha}$ affects the evolution of $f_{\alpha}$, for the $\alpha$ species. Notice that, here, we take ${\rm P}^{0}=E/a$ at order zero. We remark that in certain scenarios, such as decays where $C_{\alpha}^{(0)}\neq 0$, the first-order expansion on ${\rm P}^{0}$ from Eq. (2.3) must be included in the CN gauge [9, 13]. In the following, we assume that a possible contribution to zero order is sufficiently small so as not to spoil the background thermal form given by Eq. (2.7).
@@
In order to write the Boltzmann equation in terms of the new functions $\vartheta_{\alpha,\ell}$, we substitute (2.14) into (2.13) and those in the Fourier transformed version of equations (2.8) and (2.9). We use the comoving momentum ${\bf q}=a{\bf p}$, where ${\bf p}={\bf p}_{1}$ is the local $3$-momentum of the reference particle, and the comoving energy is determined by $\epsilon(q,\tau)=\sqrt{q^{2}+a(\tau)^{2}m^{2}}$. Then we obtain the Boltzmann hierarchy equation at first order in Fourier space:
$f_{\alpha}^{(0)}\left[\frac{\partial\vartheta_{\alpha,\ell}}{\partial\tau}+\frac{q}{\epsilon}k\left(\frac{(\ell+1)}{(2\ell+1)}\vartheta_{\alpha,\ell+1}-\frac{\ell}{(2\ell+1)}\vartheta_{\alpha,\ell-1}\right)\right]+q\frac{\partial f_{\alpha}^{(0)}}{\partial q}f_{g}=a\tilde{C}_{\ell}\,$, (2.16)
where $f_{g}$ depends on gauge choice as
$f_{g}=\frac{\partial\phi}{\partial\tau}\delta_{0\ell}+\frac{k}{3}\frac{\epsilon}{q}\psi\delta_{1\ell}$ for the CN gauge, (2.17)
$f_{g}=\frac{\dot{h}+6\dot{\eta}}{15}\delta_{2\ell}-\frac{\dot{h}}{6}\delta_{0\ell}$ $\qquad\text{for the synchronous gauge}\,$, (2.18)
@@ 2.2 Boltzmann hierarchy equations in terms of particle multipole moments
In the most general case of interacting particles, Eqs. (2.16) become a set of integro-differential ones, which are quite difficult to solve exactly [89, 97]. Since we aim to describe perturbations around a smooth universe, we introduce a perturbation to the distribution function. This perturbation is characterized by the temperature fluctuation, $\delta T_{\alpha}({\bf x},{\bf p},\tau)$, which depends explicitly on the magnitude of the momentum $p$. In contrast to the standard photon case, whose temperature anisotropies can be treated independently of the momentum magnitude due to tight coupling in the early Universe (see, for example, [50]), other particles, such as massive neutrinos, require the momentum-dependent treatment after decoupling. Then, expanding around small temperature fluctuations, $\delta T_{\alpha}$, to first order, we can write
@@
Notice that the factor of $4$ in Eq. (2.22) is a conventional choice within the standard relativistic treatment of the $\ell$-th multipole moment [39], and it also appears in the “separable ansatz” adopted for neutrinos in Refs. [98, 35]. This normalization is chosen to reproduce the energy-density–weighted average $\langle dlnf^{(0)}/d\ln q\rangle=-4$, which reflects the fact that the energy density of a relativistic species scales as $a^{-4}$.
@@ 2.3 Boltzmann hierarchy equations for massive neutrinos
$\bigg[\frac{\partial\nu_{\ell}}{\partial\tau}+\frac{q}{\epsilon}k\left(\frac{(\ell+1)}{(2\ell+1)}\nu_{\ell+1}-\frac{\ell}{(2\ell+1)}\nu_{\ell-1}\right)\bigg]-4f_{g}=-\frac{4a}{f_{\nu_{i}}^{(0)}(q)}\frac{q^{2}}{\epsilon^{2}}\left[\frac{d\ln f_{\nu_{i}}^{(0)}(q)}{d\ln q}\right]^{-1}\tilde{C}_{\ell}\,$. (2.25)
@@ 3 General Boltzmann collision term for neutrino–neutrino elastic scattering via scalar mediator
where $g_{ij}$ are the corresponding coupling constants whose labels $i$ and $j$ run over the mass eigenstates. This Lagrangian yields the same behavior as the pseudo-scalar interaction, ${\cal L}_{\rm int}^{\prime}\sim g_{ij}\varphi\overline{\nu}_{i}\gamma_{5}\nu_{j}$, when the neutrino masses are negligible, because the squared scattering amplitudes are the same [21], viable models of this type have been discussed in Refs. [12, 36, 37, 15].
However, here we will conduct the analysis as generally as possible, so that extending the results to other types of mediators or, eventually, any light self-interacting particle (as warm DM candidates; see, for instance, Refs . [118, 119]) would be straightforward.
@@ 3.1 Neutrino–neutrino interaction with different neutrino mass eigenstates
$\tilde{C}^{(1)}_{\nu_{i}}=\frac{1}{4}\frac{g^{4}T_{\nu}}{128\pi^{3}z_{1}}\frac{d\ln f^{(0)}(p_{1})}{d\ln p_{1}}\sum_{\ell=0}^{\infty}(-i)^{\ell}(2\ell+1)\nu_{\ell}{\cal P}_{\ell}(\mu)\,{\cal K}_{\ell}^{ij}(z_{1})\,.$ (3.12)
where $\upmu=\hat{\bf p}_{1}\cdot\hat{\bf k}$ (${\bf p}_{1}$ being the local momentum of the reference particle), $z_{i}=E_{i}/T_{\nu}$ and $\mu_{i}=m_{i}/T_{\nu}$ where $T_{\nu}$ is the effective temperature of background of neutrinos and
@@
The integrals in ${\cal K}_{\ell}^{ij}(z)$ give the functional coefficients:
@@
${\cal E}_{\ell}^{ij}(z_{1})$ $=\frac{1}{q_{1}}\,\int_{\mu_{j}}^{\infty}\!dz_{2}\sqrt{z_{2}^{2}-\mu_{j}^{2}}\int_{\mu_{j}}^{z_{1}+z_{2}-\mu_{i}}\!dz_{4}\,(z_{4}^{2}-\mu_{j}^{2})\,\Psi_{S,4}\,J_{4\ell}(z_{1},z_{2},z_{4})\,.$ (3.17)
@@
$\quad-\frac{1}{B_{1}^{\prime}\sqrt{-c_{z,j}^{\prime}}}\left[\alpha_{s}\left(\bar{Q}_{ij}+\frac{\bar{Q}_{ij}^{2}b_{z,j}^{\prime}}{8B_{1}^{\prime}c_{z,j}^{\prime}}\right)-\alpha_{st}D_{1z}\right]h_{6}(q_{ij}^{\prime})$ (3.32)
@@
$h_{6}(q_{ij}^{\prime})=\begin{cases}1&\text{for $\tilde{x}_{+}^{\prime}-q^{\prime}_{ij}>0$}\\ -1&\text{for $\tilde{x}_{-}^{\prime}-q^{\prime}_{ij}<0$}\\ 0&\text{elsewhere}\end{cases}\,.$ (3.34)
@@
$\frac{\partial\nu_{\ell}}{\partial\tau}$ $+\frac{q}{\epsilon}k\left[\frac{(\ell+1)}{(2\ell+1)}\nu_{\ell+1}-\frac{\ell}{(2\ell+1)}\nu_{\ell-1}\right]-4f_{g}=-\frac{a}{f_{\nu_{i}}^{(0)}(q)}\frac{g^{4}T_{\nu}}{128\pi^{3}}\left(\frac{aT_{\nu}}{\epsilon}\right)\nu_{\ell}\,{\cal C}^{ij}_{\ell}(z_{1})\,.$ (3.35)
@@
Here, $j$ denotes the (anti)neutrino mass eigenstate, in the Dirac-like case ${\rm g_{s}}=2$ refers to the spin polarization, for neutrinos or antineutrinos, and is used to determine their contribution to the energy density, which is constrained by measurements of $N_{\rm eff}$. While $g_{s}=2$ for the Majorana case.
@@ 3.2 Interactions between neutrinos of the same mass eigenstate
Consider the scattering between neutrinos in the same mass eigenstate, i.e. mass $m_{\nu}$, so that all $\alpha_{r}\to 1$ in the ${\cal T}_{v}^{N}$ expressions. If we consider $g$ as a universal coupling, we would have a non-elastic neutrino–neutrino scattering. At the moment, however, we do not consider such a scenario, since the number of particles, in principle, is not conserved in that case.
@@
${\cal D}_{\ell}(z_{1})$ $=\frac{1}{q_{1}}\,\int_{\mu_{\nu}}^{\infty}\!dz_{2}\sqrt{z_{2}^{2}-\mu_{\nu}^{2}}\int_{\mu_{\nu}}^{z_{1}+z_{2}-\mu_{\nu}}\!dz_{3}\,(z_{3}^{2}-\mu_{\nu}^{2})\,\Psi_{S,3}\,J_{3\ell}(z_{1},z_{2},z_{3})\,.$ (3.41)
@@
${\cal T}_{y}^{M}$ $=\frac{1}{\sqrt{-a_{y}}}\left(3+\frac{A^{2}}{{B_{y}^{\prime}}^{2}}-\frac{B_{y}-B_{y}^{\prime\prime}}{B_{y}^{\prime}B_{1}^{\prime}}\right)+\frac{1}{B_{1}^{\prime}\sqrt{-c_{y}^{\prime}}}\left(Q-\frac{A_{y}+B_{y}q_{y}}{B_{y}^{\prime}}-\frac{Q^{2}b_{y}^{\prime}}{8B_{1}^{\prime}c_{y}^{\prime}}\right)h_{2}(q_{y})$ (3.44)
$\quad-\frac{1}{B_{1}^{\prime}\sqrt{-c_{y}^{\prime\prime}}}\left(Q-\frac{A_{y}^{\prime\prime}+B_{y}^{\prime\prime}q^{\prime}}{B_{y}^{\prime}}+\frac{Q^{2}b_{y}^{\prime\prime}}{8B_{1}^{\prime}c_{y}^{\prime\prime}}\right)h_{3}(q^{\prime})+L_{y}(q^{\prime},q_{y})\,$,
@@
$\quad+\frac{1}{B_{1}\sqrt{-c_{x}^{\prime\prime}}}\left(C_{2x}-Q-\frac{D_{x}+B_{x}^{\prime\prime}q_{x}}{B_{x}^{\prime}}-\frac{Q^{2}b_{x}^{\prime\prime}}{8B_{1}c_{x}^{\prime\prime}}\right)h_{1}(q_{x})\,.$
@@
$A_{y}^{\prime\prime}$ $=A_{5}+B_{3}y\,,\,B_{x}^{\prime}=A_{4}^{\prime}-B_{1}^{\prime}x\,,\,B_{x}^{\prime\prime}=B_{3}-B_{1}B_{1}^{\prime}x\,,\,A_{x}^{\prime\prime}=A_{5}+B_{3}^{\prime}x$

### v2

@@ 1 Introduction
From the data analysis point of view, the main difference between these approximations relies on the cosmological scales being affected, with the mediator mass having an inverse correlation to the scale. The very light mediator and resonant regimes affect the largest scales, modifying the CMB. On the other hand, for the heavy mediator, small-scale data, especially the full shape of the matter power spectrum as seen in the galaxy and Lyman-$\alpha$ forest distributions [29, 77, 30, 106, 78] (in some scenarios) prefers a large non-null interaction coupling that should be observed in particle detectors unless one proposes or recalls a mechanism to avoid experimental bounds. Thus, there is an abnormal signal that future data may be able to explore [16, 75, 80, 81, 20, 23, 27, 15, 110, 58, 33, 28, 25, 48, 57, 113, 90, 39, 118, 114, 14, 42, 112, 61, 56, 32, 60, 69, 94, 1, 36, 86, 7, 34, 62, 54, 31, 52, 109, 64, 63, 24, 6, 95, 50, 47, 124, 44, 67, 119, 68, 2].
@@ 2.1 Boltzmann hierarchy equations under cosmological perturbations
It is clear that $Q_{\alpha}$ affects the evolution of $f_{\alpha}$, for the $\alpha$ species. Notice that, here, we take ${\rm P}^{0}=E/a$ at order zero. We remark that in certain scenarios, such as decays where $Q_{\alpha}^{(0)}\neq 0$, the first-order expansion on ${\rm P}^{0}$ from Eq. (2.3) must be included in the CN gauge [10, 14]. In the following, we assume that a possible contribution to zero order is sufficiently small so as not to spoil the background thermal form given by Eq. (2.7).
@@
In order to write the Boltzmann equation in terms of the new functions $\vartheta_{\alpha,\ell}$, we substitute (2.14) into (2.13) and those in the Fourier transformed version of equations (2.8) and (2.9). We use the comoving momentum ${\bf q}=a{\bf p}$, where ${\bf p}={\bf p}_{1}$ is the local 33-momentum of the reference particle, and the comoving energy is determined by $\epsilon(q,\tau)=\sqrt{q^{2}+a(\tau)^{2}m^{2}}$. Then we obtain the Boltzmann hierarchy equation at first order in Fourier space:
$f_{\alpha}^{(0)}\left[\frac{\partial\vartheta_{\alpha,\ell}}{\partial\tau}+\frac{q}{\epsilon}k\left(\frac{(\ell+1)}{(2\ell+1)}\vartheta_{\alpha,\ell+1}-\frac{\ell}{(2\ell+1)}\vartheta_{\alpha,\ell-1}\right)\right]+q\frac{\partial f_{\alpha}^{(0)}}{\partial q}f_{g,\ell}=a\tilde{C}_{\ell}\,$, (2.16)
where $f_{g,\ell}$ depends on gauge choice as
$f_{g,\ell}=\frac{\partial\phi}{\partial\tau}\delta_{0\ell}+\frac{k}{3}\frac{\epsilon}{q}\psi\delta_{1\ell}$ for the CN gauge, (2.17)
$f_{g,\ell}=\frac{\dot{h}+6\dot{\eta}}{15}\delta_{2\ell}-\frac{\dot{h}}{6}\delta_{0\ell}$ $\qquad\text{for the synchronous gauge}\,$, (2.18)
@@ 2.2 Boltzmann hierarchy equations in terms of particle multipole moments
In the most general case of interacting particles, Eqs. (2.16) become a set of integro-differential ones, which are quite difficult to solve exactly [91, 100]. Since we aim to describe perturbations around a smooth universe, we introduce a perturbation to the distribution function. This perturbation is characterized by the temperature fluctuation, $\delta T_{\alpha}({\bf x},{\bf p},\tau)$, which depends explicitly on the magnitude of the momentum pp. In contrast to the standard photon case, whose temperature anisotropies can be treated independently of the momentum magnitude due to tight coupling in the early Universe (see, for example, [51]), other particles, such as massive neutrinos, require the momentum-dependent treatment after decoupling. Then, expanding around small temperature fluctuations, $\delta T_{\alpha}$, to first order, we can write
@@
Notice that the factor of 44 in Eq. (2.22) is a conventional choice within the standard relativistic treatment of the $\ell$-th multipole moment [40], and it also appears in the “separable ansatz” adopted for neutrinos in Refs. [101, 36]. This normalization is chosen to reproduce the energy-density–weighted average $\langle dlnf^{(0)}/d\ln q\rangle=-4$, which reflects the fact that the energy density of a relativistic species scales as $a^{-4}$.
@@ 2.3 Boltzmann hierarchy equations for massive neutrinos
$\bigg[\frac{\partial\nu_{\ell}}{\partial\tau}+\frac{q}{\epsilon}k\left(\frac{(\ell+1)}{(2\ell+1)}\nu_{\ell+1}-\frac{\ell}{(2\ell+1)}\nu_{\ell-1}\right)\bigg]-4f_{g,\ell}=-\frac{4a}{f_{\nu_{i}}^{(0)}(q)}\frac{q^{2}}{\epsilon^{2}}\left[\frac{d\ln f_{\nu_{i}}^{(0)}(q)}{d\ln q}\right]^{-1}\tilde{C}_{\ell}\,$. (2.25)
@@ 3 General Boltzmann collision term for neutrino–neutrino elastic scattering via scalar mediator
where $g_{ij}$ are the corresponding coupling constants whose labels $i$ and $j$ run over the mass eigenstates. This Lagrangian yields the same behavior as the pseudo-scalar interaction, ${\cal L}_{\rm int}^{\prime}\sim g_{ij}\varphi\overline{\nu}_{i}\gamma_{5}\nu_{j}$, when the neutrino masses are negligible, because the squared scattering amplitudes are the same [22]. Viable models of this type have been discussed in Refs. [13, 37, 38, 16].
However, here we will conduct the analysis as generally as possible, so that extending the results to other types of mediators or, eventually, any light self-interacting particle (as warm DM candidates; see, for instance, Refs. [9, 122, 123], as well as models of self-interacting dark radiation [41]) would be straightforward.
@@ 3.1 Neutrino–neutrino interaction with different neutrino mass eigenstates
$\tilde{C}^{(1)}_{\nu_{i}}=\frac{1}{4}\frac{g^{4}T_{\nu}}{128\pi^{3}z_{1}}\frac{d\ln f^{(0)}(p_{1})}{d\ln p_{1}}\sum_{\ell=0}^{\infty}(-i)^{\ell}(2\ell+1)\nu_{\ell}{\cal P}_{\ell}(\mu)\,{\cal K}_{\ell}^{ij}(z_{1})\,$, (3.12)
where $\upmu=\hat{\bf p}_{1}\cdot\hat{\bf k}$ (${\bf p}_{1}$ being the local momentum of the reference particle). Hereafter, for further convenience, we introduce the dimensionless notation $z_{i}=E_{i}/T_{\nu}$ and $\mu_{i}=m_{i}/T_{\nu}$ where $T_{\nu}$ is the effective temperature of the neutrino background, and
@@
The functional coefficients in ${\cal K}_{\ell}^{ij}(z)$ above are given by the following integral expressions:
@@
${\cal E}_{\ell}^{ij}(z_{1})$ $=\frac{1}{q_{1}}\,\int_{\mu_{j}}^{\infty}\!dz_{2}\sqrt{z_{2}^{2}-\mu_{j}^{2}}\int_{\mu_{j}}^{z_{1}+z_{2}-\mu_{i}}\!dz_{4}\,(z_{4}^{2}-\mu_{j}^{2})\,\Psi_{S,4}\,J_{4\ell}(z_{1},z_{2},z_{4})\,$, (3.17)
@@
$\quad-\frac{1}{B_{1}^{\prime}\sqrt{-c_{z,j}^{\prime}}}\left[\alpha_{s}\left(\bar{Q}_{ij}+\frac{\bar{Q}_{ij}^{2}b_{z,j}^{\prime}}{8B_{1}^{\prime}c_{z,j}^{\prime}}\right)-\alpha_{st}D_{1z}\right]h_{6}(q_{ij}^{\prime})\,.$ (3.32)
@@
$h_{6}(q_{ij}^{\prime})=\begin{cases}1&\text{for $\tilde{x}_{+}^{\prime}-q^{\prime}_{ij}>0$}\\ -1&\text{for $\tilde{x}_{-}^{\prime}-q^{\prime}_{ij}<0$}\\ 0&\text{elsewhere}\end{cases}\,$, (3.34)
@@
$\frac{\partial\nu_{\ell}}{\partial\tau}$ $+\frac{q}{\epsilon}k\left[\frac{(\ell+1)}{(2\ell+1)}\nu_{\ell+1}-\frac{\ell}{(2\ell+1)}\nu_{\ell-1}\right]-4f_{g,\ell}=-\frac{a}{f_{\nu_{i}}^{(0)}(q)}\frac{g^{4}T_{\nu}}{128\pi^{3}}\left(\frac{aT_{\nu}}{\epsilon}\right)\nu_{\ell}\,{\cal C}^{ij}_{\ell}(z_{1})\,.$ (3.35)
@@
Here, $j$ denotes the (anti)neutrino mass eigenstate, in the Dirac-like case ${\rm g_{s}}=2$ refers to the spin polarization, for neutrinos or antineutrinos, and is used to determine their contribution to the energy density, which is constrained by measurements of $N_{\rm eff}$. While ${\rm g}_{s}=2$ for the Majorana case.
@@ 3.2 Interactions between neutrinos of the same mass eigenstate
Consider the scattering between neutrinos in the same mass eigenstate, i.e., mass $m_{\nu}$, so that all $\alpha_{r}\to 1$ in the ${\cal T}_{v}^{N}$ expressions. If we consider $g$ as a universal coupling, we would have a non-elastic neutrino–neutrino scattering. At the moment, however, we do not consider such a scenario, since the number of particles, in principle, is not conserved in that case.
@@
${\cal D}_{\ell}(z_{1})$ $=\frac{1}{q_{1}}\,\int_{\mu_{\nu}}^{\infty}\!dz_{2}\sqrt{z_{2}^{2}-\mu_{\nu}^{2}}\int_{\mu_{\nu}}^{z_{1}+z_{2}-\mu_{\nu}}\!dz_{3}\,(z_{3}^{2}-\mu_{\nu}^{2})\,\Psi_{S,3}\,J_{3\ell}(z_{1},z_{2},z_{3})\,$, (3.41)
@@
${\cal T}_{y}^{M}$ $=\frac{1}{\sqrt{-a_{y}}}\left(3+\frac{(A_{1}-B_{1}y)^{2}}{{B_{y}^{\prime}}^{2}}-\frac{B_{y}-B_{y}^{\prime\prime}}{B_{y}^{\prime}B_{1}^{\prime}}\right)-\frac{h_{3}(q^{\prime})}{B_{1}^{\prime}\sqrt{-c_{y}^{\prime\prime}}}\left(Q-\frac{A_{y}^{\prime\prime}+B_{y}^{\prime\prime}q^{\prime}}{B_{y}^{\prime}}+\frac{Q^{2}b_{y}^{\prime\prime}}{8B_{1}^{\prime}c_{y}^{\prime\prime}}\right)$ (3.44)
$\quad+\frac{1}{B_{1}^{\prime}\sqrt{-c_{y}^{\prime}}}\left(Q-\frac{A_{y}+B_{y}q_{y}}{B_{y}^{\prime}}-\frac{Q^{2}b_{y}^{\prime}}{8B_{1}^{\prime}c_{y}^{\prime}}\right)h_{2}(q_{y})+L_{y}(q^{\prime},q_{y})\,$,
@@
$\quad+\frac{1}{B_{1}\sqrt{-c_{x}^{\prime\prime}}}\left(C_{2x}-Q-\frac{D_{x}+B_{x}^{\prime\prime}q_{x}}{B_{x}^{\prime}}-\frac{Q^{2}b_{x}^{\prime\prime}}{8B_{1}c_{x}^{\prime\prime}}\right)h_{1}(q_{x})\,$,
@@
$A_{y}^{\prime\prime}$ $=A_{5}+B_{3}y\,,\,B_{x}^{\prime}=A_{4}^{\prime}-B_{1}^{\prime}x\,,\,B_{x}^{\prime\prime}=B_{3}-B_{1}B_{1}^{\prime}x\,,\,A_{x}^{\prime\prime}=A_{5}+B_{3}^{\prime}x\,$,